OpenStudy (anonymous):

prove that root 2+root 5 is irrational

Parth (parthkohli):

Let \(\sqrt{2} + \sqrt{5}\) be rational. Then\[\sqrt{2} + \sqrt{5} = \dfrac{p}{q}\tag{p,q integers and q not 0}\]\[\Rightarrow 2 + 5 + \sqrt{10} = \dfrac{p^2}{q^2}\]\[\Rightarrow \sqrt{10} = \dfrac{p^2}{q^2} - 7\]\[\Rightarrow \sqrt{10} = \dfrac{p^2 - 7q^2}{q^2}\]Now \(p^2 - 7q^2\) and \(q^2\) are integers. But \(\sqrt{10}\) cannot be written as a ratio of integers since it is irrational. This is a contradiction! And so \(\sqrt{2} + \sqrt{5}\) is irrational.

OpenStudy (preetha):

Parth, that is pretty impressive.

OpenStudy (anonymous):

Here's another solution. Note that:\[(\sqrt{2}+\sqrt{5})^2=7+\sqrt{10}\]and \[(\sqrt{2}+\sqrt{5})^4=(7+\sqrt{10})^2=59+14\sqrt{10}\]Therefore its clear that if:\[x=\sqrt{2}+\sqrt{5}\]then\[x^4-14x^2+39=0\]Using the rational roots theorem, the only possible rational solutions to that equation are:\[\pm 1, \pm 3, \pm 13, \pm 39\]However, plugging those numbers in shows that none of them are solutions. Hence if x is a solution to that polynomial, it is irrational.

Parth (parthkohli):

@Preetha Thanks, I just used conventional methods. Joe's solution is out of the box IMHO!